OpenStudy (anonymous):

Rewrite the equation below to express the cost (c) as a function of miles (m). The charge for a rental car is $20 plus $0.05 per mile. cost = $20 + $0.05m

OpenStudy (sparrow2):

what is a question?

OpenStudy (anonymous):

im guessing set the equation to m?

OpenStudy (alyssa_xo):

you're solving for m, I believe

OpenStudy (anonymous):

thanks

OpenStudy (alyssa_xo):

would you like me to verify your answer?

OpenStudy (anonymous):

f(m)=20+.05m

OpenStudy (anonymous):

that or M=c-20/.05

OpenStudy (alyssa_xo):

the 2nd one is correct. you can verify it by plugging in numbers. if the charge was $25, if \(f(m)=\dfrac{c-20}{0.05}\) \(\implies m=\dfrac{5}{0.05}=100\)

OpenStudy (anonymous):

thank you

OpenStudy (alyssa_xo):

no problem c: